What are the mole fractions of solute and solvent in 1.00 m aqueous solution?
1
Expert's answer
2020-12-14T14:47:45-0500

Solution:1.0 m (Molal) solution implies that 1 mole of the solute is dissolved in 1000 g of the solvent (water):


1)

"mole\\;fraction\\;of\\;solute= \\dfrac{1\\;mole}{1\\;mole\\;+\\; \\tfrac{1000g}{18( \\tfrac{g}{mole}\t)}\t}=\t \\dfrac{1}{56.5}\t=0.0177"

2)

"x_{solute}\;+\;x_{solvent}=1"


"x_{solvent}=1-0.0177\;=0.9823"


"Answer:"


"x_{solute}=0.0177"

"x_{solvent}=\;0.9823"